OpenStudy (anonymous):

1/4a+8=1/2a solve for a

OpenStudy (cgreenwade2000):

Is it \[\frac{ 1 }{ 2a } and \frac{ 1 }{ 4a }\] or is it \[\frac{ 1 }{ 2 }a \] and \[\frac{ 1 }{ 4 } a \]

OpenStudy (anonymous):

1/4a+8=1/2a 8=1/2a -1/4a 8=1/4a a=32